### v2

@@ 3.5 System Robustness with Majority Voting
One of the key strategies employed to enhance robustness is the use of majority voting, also known as self-consistency Wang et al. (2023b); using other strategies such as embedding-based approaches could also be possible Besta et al. (2024d). In KGoT, majority voting is implemented by querying the LLM multiple times when deciding the next step, whether to insert more data into the knowledge graph or retrieve existing data, when deciding which tool to use, and when parsing the final solution. This approach reduces the impact of single-instance errors or inconsistencies in various parts of the KGoT architecture, ensuring that the decisions made reflect the LLM’s most consistent reasoning paths.
3.6 Layered Error Containment & Management
@@
To manage LLM-generated syntax errors, KGoT includes LangChain’s JSON parsers that detect syntax issues. When a syntax error is detected, the system first attempts to correct it by adjusting the problematic syntax using different encoders, such as the “unicode escape” Python Software Foundation (2025a). If the issue persists, KGoT employs a retry mechanism (three attempts by default) that uses the LLM to rephrase the query/command and attempts to regenerate its output. If the error persists, the system logs it for further analysis, bypasses the problematic query, and continues with other iterations.
@@
To manage API & system related errors, such as the OpenAI code 500, the primary strategy employed is exponential backoff, implemented using the tenacity library Tenacity Developers (2024). Additionally, KGoT includes comprehensive logging systems as part of its error management framework. These systems track the errors encountered during system operation, providing valuable data that can be easily parsed and analyzed (e.g., snapshots of the knowledge graphs or responses from third-party APIs).
@@ 3.7 Implementation Details
Containerization with Docker and Sarus The KGoT system employs Docker Docker Inc. (2024) and Sarus Benedicic et al. (2019) for containerization, enabling a consistent and isolated runtime environment for all components. We containerize critical modules such as the KGoT controller, the Neo4j knowledge graph, and integrated tools (e.g., the Python Executor tool for safely running LLM-generated code with timeouts). Here, Docker provides a widely adopted containerization platform that guarantees consistency between development and production environments. Sarus, a specialized container platform designed for high-performance computing (HPC) environments, extends KGoT’s portability to HPC settings where Docker is typically unavailable due to security constraints. This integration allows KGoT to operate efficiently in HPC environments, leveraging their computational power.
Adaptability with LangChain The KGoT system harnesses LangChain LangChain Inc. (2024), an open-source platform specifically designed for creating and orchestrating LLM-driven applications. LangChain offers a comprehensive suite of tools and APIs that simplify the complexities of managing LLMs, including prompt engineering, tool integration, and the coordination of LLM outputs.
@@ 4 System Workflow
Figure 3: Comparison of different variants of KGoT with Hugging Face Agents and with Zero-Shot GPT-4o mini and GPT-4o. DR stands for Direct Retrieval. The used model is GPT-4o mini unless noted otherwise.
@@
The Enhance Pathway If the majority vote indicates an Enhance pathway, the next step involves determining the tools necessary for completing the Enhance operation [Uncaptioned image]. The system then orchestrates the appropriate tool calls based on the KG state [Uncaptioned image]. Once the required data from the tools is collected, the system generates the Enhance query or queries to modify the KG appropriately. Each Enhance query is executed [Uncaptioned image] and its output is validated. If an error or invalid value is returned, the system attempts to fix the query using a decoder or the LLM, retrying a specified number of times. If retries fail, the query is discarded, and the operation moves on. After processing the Enhance operation, the system increments the iteration count and continues until the KG is sufficiently expanded or the iteration limit is reached. This path ensures that the knowledge graph is enriched with relevant and accurate information, enabling the system to progress toward a solution effectively.
@@ 5 Evaluation
We now show advantages of KGoT over the state of the art. We focus on GAIA as this is currently the most comprehensive benchmark for general-purpose AI assistants, covering diverse domains such as web navigation, code execution, image reasoning, scientific QA, and multimodal tasks.
Comparison Baselines We focus on the Hugging Face (HF) Agents, the most competitive scheme in the GAIA benchmark for the hardest level 3 tasks with the GPT-4 class of models. We also compare to Zero Shot schemes where a model answers without any additional agent framework.
KGoT variants First, we vary the approach for knowledge extraction (graph queries vs. general-purpose language, cf. Section 2.3). For each option, we vary how the Solve operation is executed, by either having the LLM send a request to the backend (a Python script for NetworkX and a Cypher query for Neo4j) or by directly asking the LLM to infer the answer based on the KG, which we termed Direct Retrieval (DR). We also consider “fusion” runs, which simulate the effect from KGoT runs with both graph backends available simultaneously (or both Solve operation variants harnessed for each task). Fusion runs only incur negligible additional storage overhead because the generated KGs are small (up to several hundreds of nodes). Finally, we experiment with different tool sets. To focus on the differences coming from harnessing the KG, we reuse several utilities from AutoGen Wu et al. (2024) such as Browser and MDConverter, and tools from HF Agents, such as Surfer Agent, web browsing tools, and Text Inspector.
@@
Considered Dataset We use the GAIA benchmark Mialon et al. (2024); focusing on the validation set (165 tasks) for budgetary reasons and also because it comes with the ground truth answers.
Scalability We verified that selecting Neo4j as the graph query backend is not the bottleneck (the majority of the time is spent on the tool usage, most importantly, web browsing and text parsing). Moreover, due to the effective knowledge extraction process and the nature of the tasks considered (i.e., AI assistance), none of the tasks require large KGs. The maximum graph size that we observed was 522 nodes. This is orders of magnitude below any scalability concerns.
@@ 5.1 Advantages of KGoT
Figure 3 shows the number of solved tasks (left side) as well as the average cost per solved task (right side) for different KGoT variants (explained above) as well as three baselines (HF Agents using both GPT-4o mini and GPT-4o, and the Zero-Shot GPT-4o mini and GPT-4o). Additionally, we show the Pareto front in Figure 4 for the multidimensional optimization problem of improving accuracy (i.e., reducing failed tasks) and lowering cost. All variants of KGoT solve a greater number of tasks (up to 9 more) compared to HF Agents while also being more cost-efficient (between 42% to 62% lower costs). The key reason for the KGoT advantages stems from harnessing the knowledge graph–based representation of the evolving task state.
The ideal fusion runs of Neo4j and NetworkX solve an even greater number of tasks (57 for both) than the single runs, and also have a lower average cost (up to 68% lower than Hugging Face Agents). The fusion of all combinations of backend and solver types would solve by far the highest number of tasks (71) – more than twice as much as Hugging Face Agents – while also exhibiting the lowest average cost per solved query (nearly 72% lower than Hugging Face Agents).
The direct zero-shot use of GPT-4o mini and GPT-4o has the lowest average cost per solved task (just \$0.0013 and \$0.0164 respectively), making it the most cost-effective, however this approach is only able to solve 17 and 29 tasks, respectively.
Figure 4: Pareto front plot of cost and error counts. We report results for answering 165 GAIA validation questions across different comparison targets, using the GPT-4o mini model with each baseline. For the Zero-Shot inference, we also include results for GPT-4o for comparison. DR means Direct Retrieval.

### v6

@@ 3.5 Ensuring System Robustness
Robustness is ensured with two established mechanisms, Self-Consistency (Wang et al., 2023b) (via majority voting) and LLM-as-a-Judge (Gu et al., 2025) (other strategies such as embedding-based stability are also applicable (Besta et al., 2025d)). With Self-Consistency, we query the LLM multiple times when deciding whether to insert more data into the KG or retrieve existing data, when deciding which tool to use, and when parsing the final solution. This approach reduces the impact of single-instance errors or inconsistencies in various parts of the KGoT architecture. LLM-as-a-Judge further reinforces the robustness, by directly employing the LLM agent to make these decisions based on generated reasoning chains.
Overall, both Self-Consistency and LLM-as-a-Judge have been shown to significantly enhance the robustness of prompting. For example, MT-Bench and Chatbot Arena show that strong judges (e.g., GPT-4 class) match human preferences at 80% agreement or more, on par with human-human agreement (Zheng et al., 2023). Prometheus/Prometheus-2 further demonstrate open evaluator LMs with the highest correlations to both humans and proprietary judges across direct-assessment and pairwise settings, and AlpacaEval has been validated against approximately 20K human annotations, addressing earlier concerns about reproducibility at scale. Similarly reliable gains have been shown for Self-Consistency (Wang et al., 2023b).
3.6 Ensuring Layered Error Containment & Management
@@
To manage LLM-generated syntax errors, KGoT includes LangChain’s JSON parsers that detect syntax issues. When a syntax error is detected, the system first attempts to correct it by adjusting the problematic syntax using different encoders, such as the “unicode escape” (Python Software Foundation, 2025a). If the issue persists, KGoT employs a retry mechanism (three attempts by default) that uses the LLM to rephrase the query/command and attempts to regenerate its output. If the error persists, the system logs it for further analysis, bypasses the problematic query, and continues with other iterations.
@@
To handle API & system related errors, such as the OpenAI code 500, we employ exponential backoff, implemented using the tenacity library (Tenacity Developers, 2025a). Additionally, KGoT includes comprehensive logging systems as part of its error management framework. These systems track the errors encountered during system operation, providing valuable data that can be easily parsed and analyzed (e.g., snapshots of the knowledge graphs or responses from third-party APIs).
@@ 3.7 Implementation Details
KGoT employs Docker (Docker Inc., 2025) and Sarus (Benedicic et al., 2019) for containerization, enabling a consistent and isolated runtime environment for all components. We containerize critical modules such as the KGoT controller, the Neo4j knowledge graph, and integrated tools (e.g., the Python Executor tool for safely running LLM-generated code with timeouts). Here, Docker provides a widely adopted containerization platform for local and cloud deployments that guarantees consistency between development and production environments. Sarus, a specialized container platform designed for high-performance computing (HPC) environments, extends KGoT’s portability to HPC settings where Docker is typically unavailable due to security constraints. This integration allows KGoT to operate efficiently in HPC environments, leveraging their computational power.
KGoT also harnesses LangChain (LangChain Inc., 2025a), an open-source framework specifically designed for creating and orchestrating LLM-driven applications. LangChain offers a comprehensive suite of tools and APIs that simplify the complexities of managing LLMs, including prompt engineering, tool integration, and the coordination of LLM outputs.
@@ 4 System Workflow
The Enhance Pathway If the majority vote indicates an Enhance pathway, the next step involves determining the tools necessary for completing the Enhance operation [Uncaptioned image]. The system then orchestrates the appropriate tool calls based on the KG state [Uncaptioned image]. Once the required data from the tools is collected, the system generates the Enhance query or queries to modify the KG appropriately. Each Enhance query is executed [Uncaptioned image] and its output is validated. If an error or invalid value is returned, the system attempts to fix the query, retrying a specified number of times. If retries fail, the query is discarded, and the operation moves on. After processing the Enhance operation, the system increments the iteration count and continues until the KG is sufficiently expanded or the iteration limit is reached. This path ensures that the knowledge graph is enriched with relevant and accurate information, enabling the system to progress toward a solution effectively.
@@ 5 Evaluation
We now show advantages of KGoT over the state of the art. Additional results and full details on the evaluation setup are in Appendix D.
Comparison Baselines. We focus on the Hugging Face (HF) Agents (Roucher & Petrov, 2025), the most competitive scheme in the GAIA benchmark for the hardest level 3 tasks with the GPT-4 class of models. We also compare to two agentic frameworks, namely GPTSwarm (Zhuge et al., 2024) (a representative graph-enhanced multi-agent scheme) and Magentic-One (Fourney et al., 2024), an AI agent equipped with a central orchestrator and multiple integrated tool agents. Next, to evaluate whether database search outperforms graph-based knowledge extraction, we also consider two retrieval-augmented generation (RAG) (Lewis et al., 2020) schemes, a simple RAG scheme and GraphRAG (Edge et al., 2025). Both RAG baselines use the same tool-generated knowledge, chunking data at tool-call granularity (i.e., a chunk corresponds to individual tool call output). Simple RAG constructs a vector database from these tool outputs while GraphRAG instead models the tool outputs as a static KG of entities and relations, enabling retrieval via graph traversal. Finally, we use Zero-Shot schemes where a model answers without any additional agent framework.
KGoT variants. First, we experiment with graph query languages vs. general-purpose languages, cf. Section 2.3. For each option, we vary how the Solve operation is executed, by either having the LLM send a request to the backend (a Python script for NetworkX and a Cypher/SPARQL query for Neo4j/RDF4J) or by directly asking the LLM to infer the answer based on the KG (Direct Retrieval (DR)). We experiment with different query languages (Cypher vs. SPARQL). We also consider “fusion” runs, which simulate the effect from KGoT runs with both graph backends available simultaneously (or both Solve operation variants harnessed for each task). Fusion runs only incur negligible additional storage overhead because the generated KGs are small (up to several hundreds of nodes). Finally, we experiment with different tool sets. To focus on the differences coming from harnessing the KG, we reuse several utilities from AutoGen (Wu et al., 2024) such as Browser and MDConverter, and tools from HF Agents, such as Surfer Agent, web browsing tools, and Text Inspector.
@@
Considered Datasets We use the GAIA benchmark (Mialon et al., 2024) focusing on the validation set (165 tasks) for budgetary reasons and also because it comes with the ground truth answers. The considered tasks are highly diverse in nature; many require parsing websites or analyzing PDF, image, and audio files. We focus on GAIA as this is currently the most comprehensive benchmark for general-purpose AI assistants, covering diverse domains such as web navigation, code execution, image reasoning, scientific QA, and multimodal tasks. We further evaluate on SimpleQA (Wei et al., 2024), a factuality benchmark of 4,326 questions, of which we sample 10% for budgetary reasons. The dataset spans diverse topics and emphasizes single, verifiable answers, making it effective for assessing factual accuracy.
Figure 3: Advantages of different variants of KGoT over other baselines (Hugging Face Agents using both GPT-4o-mini and GPT-4o, Magentic-One, GPTSwarm, two RAG baselines, Zero-Shot GPT-4o mini, and Zero-Shot GPT-4o) on the validation dataset of the GAIA benchmark. DR stands for Direct Retrieval. The used model is GPT-4o mini unless noted otherwise.
@@ 5.1 Advantages of KGoT
Figure 3 shows the number of solved tasks (the left side) as well as the average cost per solved task (the right side) for different KGoT variants as well as all comparison baselines. While we focus on GPT-4o mini, we also show the results for HF Agents and Zero-Shot with GPT-4o. Additionally, we show the Pareto front in Figure 11 for the multidimensional optimization problem of improving accuracy (i.e., reducing failed tasks) and lowering cost. All variants of KGoT solve a greater number of tasks (up to 9 more) compared to HF Agents while also being more cost-efficient (between 42% to 62% lower costs). The key reason for the KGoT advantages stems from harnessing the knowledge graph–based representation of the evolving task state.
The ideal fusion runs of Neo4j and NetworkX solve an even greater number of tasks (57 for both) than the single runs, they have a lower average cost (up to 62% lower than HF Agents), and they even outperform HF Agents with GPT-4o. The fusion of all combinations of backend and solver types solve by far the highest number of tasks (71) – more than twice as much as HF Agents – while also exhibiting 44% lower cost than HF Agents. The direct Zero-Shot use of GPT-4o mini and GPT-4o has the lowest average cost per solved task (just \$0.0013 and \$0.0164 respectively), making it the most cost-effective, however this approach is only able to solve 17 and 29 tasks, respectively. GPTSwarm is cheaper compared to KGoT, but also comes with fewer solved tasks (only 26). While Magentic-One is a capable agent with a sophisticated architecture, its performance with GPT-4o mini is limited, solving 31 tasks correctly, while also exhibiting significantly higher costs. Simple RAG yields somewhat higher costs than KGoT and it solves fewer tasks (35). GraphRAG performs even worse, solving only 23 tasks and incurring even higher cost. While neither RAG baseline can invoke new tools to gather missing information (reducing accuracy and adaptability), GraphRAG’s worse performance is due to the fact that it primarily targets query summarization and not tasks as diverse as those tested by GAIA. Overall, KGoT achieves the best cost-accuracy tradeoff, being both highly affordable and very effective.
5.2 Analysis of Methods for Knowledge Extraction
We explore different methods of extracting knowledge. Overall, in many situations, different methods have complementary strengths and weaknesses.
Graph queries with Neo4j excel at queries such as counting patterns. Yet, Cypher queries can be difficult to generate correctly, especially for graphs with more nodes and edges. Despite this, KGoT’s Cypher queries are able to solve many new GAIA tasks that could not be solved without harnessing Cypher. SPARQL (Pérez et al., 2009) + RDF4J (Eclipse Foundation, 2025) is slightly worse (36 tasks solved) than Cypher + Neo4j (existing literature also indicates that LLMs have difficulties formulating effective SPARQL queries (Emonet et al., 2024; Mecharnia & d’Aquin, 2025)).
Python with NetworkX offers certain advantages over Neo4j by eliminating the need for a separate database server, making it a lightweight choice for the KG. Moreover, NetworkX computations are fast and efficient for small to medium-sized graphs without the overhead of database transactions. Unlike Neo4j, which requires writing Cypher queries, we observe that in cases where Neo4j-based implementations struggle, NetworkX-generated graphs tend to be more detailed and provide richer vertex properties and relationships. This is likely due to the greater flexibility of Python code over Cypher queries for graph insertion, enabling more fine-grained control over vertex attributes and relationships. Another reason may be the fact that Python is likely more represented in the training data of the respective models than Cypher.
Our analysis of failed tasks indicates that, in many cases, the KG contains the required data, but the graph query fails to extract it. In such scenarios, Direct Retrieval, where the entire KG is included in the model’s context, performs significantly better by bypassing query composition issues. However, Direct Retrieval demonstrates lower accuracy in cases requiring structured, multi-step reasoning.
5.3 Advantages on the GAIA Test Set
Table 1: Comparison of KGoT with other current state-of-the-art open-source agents on the full GAIA test set. The baseline data, including for TapeAgent (Bahdanau et al., 2024), of the number of solved tasks is obtained through the GAIA Leaderboard (Mialon et al., 2025). We highlight the best performing scheme in a given category in bold. Model: GPT-4o mini.
Agents All L1 L2 L3
GPTSwarm 33 15 15 3
Magentic-One 43 22 18 3
TapeAgent 66 28 35 3
Hugging Face Agents 68 30 34 4
KGoT (fusion) 73 33 36 4
Furthermore, our approach achieves state-of-the-art performance on the GAIA test set with the GPT-4o mini model. The results are shown in Table 1, underscoring its effectiveness across all evaluation levels. The test set consists of 301 tasks (93 level 1 tasks, 159 level 2 tasks and 49 level 3 tasks).
5.4 Advantages beyond GAIA Benchmark
We also evaluate KGoT as well as HF Agents and GPTSwarm on a 10% sample (433 tasks) of the SimpleQA benchmark (detailed results are in Appendix D.1). KGoT performs best, solving 73.21%, while HF Agents and GPTSwarm exhibit reduced accuracy (66.05% and 53.81% respectively). KGoT incurs only 0.018\$ per solved task, less than a third of the HF Agents costs (0.058\$), while being somewhat more expensive than GPTSwarm (0.00093\$).
[…]
5.5 Ensuring Scalability and Mitigating Bottlenecks
The primary bottleneck in KGoT arises from I/O-bound and latency-sensitive LLM tool invocations (e.g., web browsing, text parsing), which account for 72% of the runtime, which KGoT mitigates through asynchronous execution and graph operation parallelism as discussed in Section 3.4. A detailed breakdown of the runtime is reported in Appendix D.3. Figure 10 confirms KGoT’s scalability, as increasing the number of parallelism consistently reduces the runtime. Moreover, due to the effective knowledge extraction process and the nature of the tasks considered, none of the tasks require large KGs. The maximum graph size that we observed was 522 nodes. This is orders of magnitude below any scalability concerns.